OpenStudy (anonymous):

9x^2+4y^2+36x-24y+36=0 find vertex foci and center

terenzreignz (terenzreignz):

This is an ellipse, no?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Okay... group terms together, x's together, y's together \[\large \color{red}{9x^2+36x}+\color{blue}{4y^2-24y}+36 = 0\]

terenzreignz (terenzreignz):

\[\large \color{red}{9(x^2+4x)}+\color{blue}{4(y^2-6y)}+36 = 0\]

OpenStudy (anonymous):

ok

terenzreignz (terenzreignz):

You know how to complete the square?

OpenStudy (anonymous):

yea

terenzreignz (terenzreignz):

So, I want you to complete the square of the red (x) part

OpenStudy (anonymous):

so you add 4 to 36 and 4 to 0

OpenStudy (anonymous):

and multiply the 4 by 9

terenzreignz (terenzreignz):

\[\large \color{red}{9(x^2+4x+4)}+\color{blue}{4(y^2-6y)}+36 = 0+\color{red}{(4\times9)}\]

terenzreignz (terenzreignz):

You add the 4 within the parenthesis.

OpenStudy (anonymous):

oh ok

terenzreignz (terenzreignz):

what to add the blue (y) part?

OpenStudy (anonymous):

same thing but you add3 times 4

terenzreignz (terenzreignz):

you don't add 3

terenzreignz (terenzreignz):

complete the square, remember?

OpenStudy (anonymous):

ya i know after completing the square i got 3 so i add 3 within the parantheses and 3(9) after the equal sign

terenzreignz (terenzreignz):

you take half of 6 and then SQUARE it.

OpenStudy (anonymous):

ooops i just added wrong

terenzreignz (terenzreignz):

so... fix it :)

OpenStudy (anonymous):

ok its 9

OpenStudy (anonymous):

give me a minute to write my equation

terenzreignz (terenzreignz):

sure.

OpenStudy (anonymous):

9(x^{2}+4x+4) + 4(y^{2}-6y+9) +36 = 9(4)+ 4(9)

OpenStudy (anonymous):

= 72

terenzreignz (terenzreignz):

Nicely done :) \[\large \color{red}{9(x^2+4x+4)}+\color{blue}{4(y^2-6y)}+36 = 0+\color{red}{(4\times9)}+\color{blue}{(9\times 4)}\] \[\large \color{red}{9(x^2+4x+4)}+\color{blue}{4(y^2-6y+9)}+36 = \color{violet}{72}\]

OpenStudy (anonymous):

9(x+2)^2 + 4 ( x-3 ) ^ 2 + 36 = 72

OpenStudy (anonymous):

now divide by 72 on both sides

terenzreignz (terenzreignz):

You're... good at this :D \[\large \color{red}{9(x+2)^2}+\color{blue}{4(y-3)^2}+36 = 72\]

terenzreignz (terenzreignz):

Not yet, you still have a pesky 36 on the left side :)

OpenStudy (anonymous):

oh so you move it to the other side

terenzreignz (terenzreignz):

yup :)

OpenStudy (anonymous):

= 108

OpenStudy (anonymous):

divide by 108 on both sides

terenzreignz (terenzreignz):

I don't think so... you subtract 36 on the right side, don'cha? :)

OpenStudy (anonymous):

ugh stupid mistakes

terenzreignz (terenzreignz):

Don't worry, that's why I'm here :)

OpenStudy (anonymous):

= 36

OpenStudy (anonymous):

( x+2) ^2 div 4 + ( x-3)^2 div 9 = 1

OpenStudy (anonymous):

now a = 3 and b = 2 so its a vertical ellipse

terenzreignz (terenzreignz):

Yeah, it would seem so. :) and no, it's a horizontal ellipse if that's the case :D

OpenStudy (anonymous):

no its vertical

terenzreignz (terenzreignz):

oh yes... it's b that's 3 and a that's 2...

OpenStudy (anonymous):

so its vertical ?

terenzreignz (terenzreignz):

yup

OpenStudy (anonymous):

center = (-2,3)

terenzreignz (terenzreignz):

\[\large \color{red}{\frac{(x+2)^2}{2^2}}+\color{blue}{\frac{(y-3)^2}{3^2}} = 1\]

OpenStudy (anonymous):

yes i o put x on accident before

OpenStudy (anonymous):

center = (-2,3)

terenzreignz (terenzreignz):

no worries. correct with your centre. Now your vertices?

OpenStudy (anonymous):

(-2,6) and (-2,5)

terenzreignz (terenzreignz):

nope...

OpenStudy (anonymous):

c= \[2\sqrt{3}\]

OpenStudy (anonymous):

why not doesnt vertex = ( h,k+a)

terenzreignz (terenzreignz):

Sure, also (h,k-a)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so ( -2,6) and (-2, 5 )

terenzreignz (terenzreignz):

How do you get (-2,5)?

OpenStudy (anonymous):

oooh its supposed to be (-2,0)

terenzreignz (terenzreignz):

much better. Now your foci? You got what value for c?

OpenStudy (anonymous):

\[2\sqrt{3}\]

terenzreignz (terenzreignz):

and so your foci are at...?

terenzreignz (terenzreignz):

Wait, no, I don't think so...

OpenStudy (anonymous):

( 3,3+ c ) and

terenzreignz (terenzreignz):

\[c = \sqrt{a^2-b^2}\]

OpenStudy (anonymous):

oh ya its square nroot of 13 not 12 so c is equal to 3.6

terenzreignz (terenzreignz):

c must not be bigger than a. Or else your foci would be outside the ellipse. I think you calculated \[\huge \sqrt{a^2 + b^2}\] when in fact \[\huge \color{blue}{c=\sqrt{a^2-b^2}}\]

OpenStudy (anonymous):

ok ok i got this foci = ( 3, 5.25 ) and ( 3, .75)

terenzreignz (terenzreignz):

what did you get for c?

OpenStudy (anonymous):

2.25

terenzreignz (terenzreignz):

how? You're actually supposed to get something like 2.236

OpenStudy (anonymous):

your right im blind

terenzreignz (terenzreignz):

okay, well, as long as you got c correctly, your foci should be at (h , k + c) and (h , k - c) Do that, and you're done :)

OpenStudy (anonymous):

ok now he gave us another question similiar so im gonna try doing that then come back to review just the answers

OpenStudy (anonymous):

thank you so much for your time

terenzreignz (terenzreignz):

You're welcome :)

OpenStudy (anonymous):

are you still there ?